OpenStudy (anonymous):

Two forces act on a 1250 kg sailboat as it moves through the water with an initial velocity of 11 m/s. The forward force of the wind exerts 3.90*10^3 N while the resistive drag force of the water exerts 4.35 * 10^3 N. What's the velocity of the boat 15 seconds later?

OpenStudy (anonymous):

resultant force on the boat = 3.9 *10^3 - 4.35*10^3 = - 0.45* 10^3 N so the boat is decelerating -0.45* 10^3 = 1250 * a (newtons 2nd law) a = -0.45*10^3 / 1250 = -0.36 m/s/s using v = u + at = 11 - 0.36*15 = 5.6 m/s

OpenStudy (anonymous):

OHMEHGLOB! THANK YOU SO MUCH!!